# Virtual Width Networks

### v1

@@ 1 Introduction
We evaluate VWN across multiple regimes using internal MoE models. We report training dynamics and token efficiency relative to matched non‑VWN baselines, and assess downstream generalization. Headline results show that VWN, which expands the embedding width by 8×, achieves the baseline’s next‑token loss with $2.5\times$ fewer tokens and the next‑2‑token loss with $3.5\times$ fewer tokens, with the efficiency advantage increasing as training progresses, as shown in Fig. 1.
@@ 3.2 Over-Width Embedding
To increase the embedding dimensions, we propose the Over-Width Embedding technique. Given a fixed hidden size $D$, we enlarge the embedding dimension to a wider dimension $D^{\prime}$, resulting in richer token embeddings without a substantial increase in computational overhead.
@@ 3.4 Multi-token Prediction
As for the output layer, previous research [huang2025over] has demonstrated that Multi-Token Prediction (MTP) serves as an approximation of $k$-gram decoding. Building upon this insight, we leverage MTP to provide fine-grained supervised signals by introducing additional VWN layers atop the backbone model, thus constructing an enhanced prediction head. Specifically, following deepseekai2025deepseekv3technicalreport, we concatenate the embedding of the next token with the last-layer embedding of the preceding token, applying a linear projection to generate logits, as illustrated in the upper part of Fig 3 (c).
@@ 5.2 Large Virtual Width
The observed loss reductions follow a log‑linear relation with respect to the virtual‑width factor $r$ (Fig. 8). A fitted coefficient of 0.0069-0.0069 indicates that each doubling of virtual width corresponds to an approximate loss reduction of 0.0069. While the effect size is modest, it suggests a systematic efficiency gain attributable to virtual widening. We hypothesize that more expressive backbones and improved mechanisms that more effectively leverage the virtual‑width hidden representations could further amplify the efficiency gains observed with VWN.

### v2

@@ 1 Introduction
We evaluate VWN across multiple regimes using internal MoE models. We report training dynamics and token efficiency relative to matched non‑VWN baselines, and assess downstream generalization. Headline results show that VWN, which expands the embedding width by 8×, achieves the baseline’s next‑token loss with $2.5\times$ fewer tokens and the next‑2‑token loss with $3.5\times$ fewer tokens, with the efficiency advantage increasing as training progresses, as shown in Figure 1.
@@ 3.2 Over-Width Embedding
To increase the embedding dimensions, we propose the Over-Width Embedding technique. Given a fixed hidden size $D$, we enlarge the embedding dimension of input to a wider dimension $D^{\prime}$, resulting in richer token embeddings without a substantial increase in computational overhead.
@@
For unembedding, the model needs to map the last over‑width hidden states back to the original hidden width $D$ before the unembedding layer. We introduce a reduce operator $\mathbf{W}_{\text{reduce}}\in\mathbb{R}^{D\times D^{\prime}}$ that performs a linear projection from the over‑width dimension $D^{\prime}$ to the original width $D$:
$\mathbf{h}^{L}_{\text{reduce}}=\mathbf{W}_{\text{reduce}}\,\mathbf{h}^{\prime L}.$ (4)
[…]
@@ 3.4 Multi-token Prediction
As for the output layer, previous research [huang2025over] has demonstrated that Multi-Token Prediction (MTP) serves as an approximation of $k$-gram decoding. Building upon this insight, we leverage MTP to provide fine-grained supervised signals by introducing additional VWN layers atop the backbone model, thus constructing an enhanced prediction head. Specifically, following deepseekai2025deepseekv3technicalreport, we concatenate the embedding of the next token with the last-layer embedding of the preceding token, applying a linear projection to generate logits, as illustrated in the upper part of part of Figure 3 (c).
@@ 5.2 Large Virtual Width
The observed loss reductions follow a log‑linear relation with respect to the virtual‑width factor $r$ (Figure 8). A fitted coefficient of 0.0069-0.0069 indicates that each doubling of virtual width corresponds to an approximate loss reduction of 0.0069. While the effect size is modest, it suggests a systematic efficiency gain attributable to virtual widening. We hypothesize that more expressive backbones and improved mechanisms that more effectively leverage the virtual‑width hidden representations could further amplify the efficiency gains observed with VWN.